OpenStudy (anonymous):

2^i how to i simplify

OpenStudy (solomonzelman):

2^i is a very weird number. Just like saying that you have 3½ friends.... if you know what I mean, you are multiplying 2 "i" number of times, even though, "i" doesn't really have a value. I can simplify i^2, but if it is 2^i, I would not be playing around with it.

OpenStudy (anonymous):

\[i = \sqrt{-1}\] \[2^{i} = 2^{\sqrt{-1}}\] That's all I can come up with.

OpenStudy (anonymous):

\[\Large 2^i=\left(e^{\ln2}\right)^i=e^{i\ln2}=\cos(\ln2)+i\sin(\ln2)\]

OpenStudy (solomonzelman):

and that is to "simplify" ?

OpenStudy (anonymous):

Well I would think a complex number in \(a+bi\) form is easier to work with, no?